OpenStudy (anonymous):

help it the last question plz Use the rectangle below. Show all work. a. Find the perimeter. b. Find the area.

OpenStudy (anonymous):



OpenStudy (anonymous):

the other edge is square root of 15^2 - 12^2=9^2 so it's 9, area is 9*12 = 108 cm

OpenStudy (anonymous):

thx